News & Status
  • News & status not loaded yet.

Javascript is turned off in your web browser.

To take full advantage of this site, please turn it on and then refresh the page.

Microscope Facility News & Equipment Status

To change News view:
Two months
Quarter
Year
All
Select multiple with control-key
Restrict listing
to selected
equipment
(Exclude announcements)

2015-05-11 (Zeiss LSM 780 FCS): 780 Argon Laser (CFP, GFP, YFP) - REDUCED POWER

Update (5/12, 11:34am) The 780 argon laser has been replaced and Barbara is running some tests right now. It will be ready for the 12pm reservation today. Thank you all.

Update (5/11, 12:50pm) The new Argon laser has arrived and the engineer will install it Tuesday morning. Loza

The Zeiss LSM 780 has a reduced Argon laser power – this is the power for the 458, 488 and 514 laser lines (i.e. CFP, GFP, YFP). We have placed a service call with Zeiss but we want to inform users that while the scope is still operational, you will need to use a higher laser power and/or gain to image your samples on those channels/lines. I was able to take images on our standard control slides, but if you are unable to image on the system, please note this in the logbook. If you wish to cancel your reservation(s) and are unable to do so, please e-mail microscopy@jhmi.edu. We apologize for any inconvenience and will send further updates as they become available.

Thank you,
Barbara

2015-04-24 (Leica Cryostat): Routine Cryostat Maintenance

Hi Cryostat Users, Starting 2/24/15 at 4pm and ending on Monday 2/27/15 at 1pm the cryostat will be unavailable for use due to routine defrosting and cleaning. I apologize for any inconvenience this may cause. --Tyler

2015-04-14 (Philips CM120 TEM): Philips TEM: serviced on 4/9-4/10

We have been experiencing vacuum loss when the sample holder is removed and replaced. The microscope was serviced by FEI on 4/9-4/10/2015. During the engineer's visit, the airlock was cleaned, the o-ring on holder was replaced and scratches on the holder were polished out. The engineer also did a beam check and adjusted the TEM for 80kv operation. Normal operations resumed on 4/10/2015. (Source: FEI Field Service Report KEW01097)

Last modified